How many grams of Lactic acid (C6H10O5) should be dissolved in 650g of cyclohexane (C6H12) to raise the boiling point to 84.910C? The normal boiling point of cyclohexane is 810C and its molal boiling point is 2.75 0C/m.

1
Expert's answer
2022-02-19T06:55:37-0500

The formula of lactic acid is not C6H10O5, it is C3H6O3. Moreover, it is unlikely to be well soluble in cyclohexane. Anyway, the solution below will be provided for the formula as given in the task (C6H10O5).


"\Delta{T}=k_b\times{m}" , where m is the molality. Solving for m,


"m=\frac{\Delta{T}}{k_b}=\frac{84.91\degree{C}-81\degree{C}}{2.75\degree{C}/m}=1.42\frac{mol\ C_6H_{10}O_5}{kg\ C_6H_{12}}"


The molar mass of C6H10O5 is 162.14 g/mol


"mass(C_6H_{10}O_5)=0.650\ kg(C_6H_{12})\times\frac{1.42\ mol\ C_6H_{10}O_5}{1\ kg\ C_6H_{12}}\times\frac{162.14\ g\ C_6H_{10}O_5}{1\ mol\ C_6H_{10}O_5}=150\ g"


Answer: 150 g